A 5 kg Blocks slides along a rough horizontal surface the coefficient of kinetic friction for the block and the surface is 0.4 at what rate in watts is the frictional force doing work on the block at the instant when it’s Speed s 5 m/s
1
Expert's answer
2020-08-10T19:53:50-0400

The power

"P=\frac{dW}{dt}=Fv=\mu mgv"

"P=0.4\times (5\:\rm kg)\times (9.8\: N/kg)\times (5\: m/s)=98\: W"